[Brahms-l] Shift report 20050204 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 04 2005 - 00:11:56 EST
1. Beam delivered from 17.30

2. MRS@45 B2255, FS@12 B1219
run    tr2   tr3
12762  33k  152k
12763  34k  233k
12764  29k  204k
12765  15k  195

MRS@45 zero field, FS@12 zero field (D1@B430)
12766   3k   44k
(This was taken the last 10 min of the store...)

3. None seen, except that MRS was at 45 not 40 deg. This has been changed in the Database by Hiro for all runs since the Power outage

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 4 00:12:21 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 04 2005 - 00:12:44 EST